# Two competing populations with a common environmental resource thanks:

### v1

@@ I Introduction
Classical game theory describes strategic interactions between individuals provided that the incentives for their choices are static, i.e. do not change over time. However, their choices often have an impact on a shared environment, which in turn may affect their incentives for future decisions. The utilization of common resources such as water, fishing grounds, or traffic networks best illustrates this – high individual utilization makes fewer resources available to others in the future [1]. Clearly, a dynamic interplay exists between users’ decisions and their impact on a shared environmental state.
@@
Much of the existing research focuses on a single population of agents whose decisions impact the environment in isolation [8, 9, 10, 7, 11, 12, 13]. These models are not sufficient to describe multi-population interactions. For example, individuals in neighboring countries follow different environmental policies, yet utilize resources from the same common source (e.g. fish in the ocean, water from rivers, clean air). Extensions of feedback-evolving games to multiple populations, and possibly multiple local environments, would enable a vastly richer set of scenarios for study. While some work has studied the dynamics of multi-population models in recent years [14, 15, 16], an interesting and understudied direction involves hierarchical decision-making, i.e. the strategic selection of local incentive policies, under the context that the activity from one population has externalities on resources utilized by other populations.
@@ II-A Preliminary on Feedback-evolving games
On the other hand, the $A_{0}$ matrix describes payoffs when the environment is depleted. We interpret $A_{0}$ to be an “environmental policy” that the population has follows. For example, the low consumption strategy may be more incentivized when the environment is bad (e.g. subsidies for using electric vehicles). The payoff structure of the $A_{0}$ matrix is completely determined from the parameters $\delta_{SP0}:=S_{0}-P_{0}$ and $\delta_{RT0}:=R_{0}-T_{0}$.
@@
The state of the system $(x,n)$ evolves over the state space $[0,1]^{2}$. Indeed, the set $(0,1)^{2}$ is forward-invariant under the dynamics (3). The originating work [2] provided a full characterization of the asymptotic outcomes of system (3) for all possible environmental policy matrices $A_{0}$.
@@
2) Oscillating Tragedy of the commons: If $\delta_{SP0}>0$ and $\frac{\delta_{TR1}}{\delta_{PS1}}\delta_{SP0}<\delta_{RT0}$, then system (3) exhibits a stable heteroclinic cycle between the four corner fixed points, $(0,0)$, $(1,0)$, $(1,1)$, $(0,1)$.
3) Tragedy of the commons: If $\delta_{SP0}\leq 0$, or $\delta_{SP0}>0$ and $\delta_{RT0}<-\frac{\theta}{\alpha}\delta_{SP0}$, then the only asymptotically stable fixed point has $n=0$.
The environmental policies belonging to $\mathcal{V}$ are ones that can sustain a stable and non-zero resource level. In other words, they avert a tragedy of the commons [17], which refers to a fixed point where the resource level is zero. The policies in $\mathcal{V}$ can be considered as more desirable operating conditions than the policies described in items 2 and 3 above. The oscillating tragedy of the commons is considered an undesirable outcome, in which the system cycles between states of nearly zero resource levels, and states of fully abundant resource levels. Visually, the state trajectory “spirals outward” from the interior, and its $\omega$-limit set is the boundary of the state space.
@@ II-B Model: Two-population feedback-evolving games
where $h(x_{1},x_{2}):=\sum_{i=1}^{2}(\theta_{i}x_{i}-\alpha_{i}(1-x_{i}))$, and with initial condition $(x_{1}(0),x_{2}(0),n(0))\in(0,1)^{3}$. By construction, $(0,1)^{3}$ is a forward-invariant set. We are interested in studying how the dynamics of the two-population environmental coupled system (7) behaves with respect to the population’s defining parameters – especially how the environmental policies $A_{0}^{(1)}$, $A_{0}^{(2)}$ adopted by each population affects environmental resources. We will make the following assumptions on these environmental policies.
@@
Assumption 2 asserts that the relative payoff to low consumers in both populations monotonically decreases as the environmental state improves. The next and final assumption focuses our study on a scenario where one population is “environmentally responsible”, and the other population is “irresponsible”.
@@
The first part of Assumption 3 asserts that agents in population 1 are cooperative enough to sustain the resource (Theorem 2.1, item 1) in the absence of population 2. The second part asserts that agents in population 2 make no effort to conserve resources, always preferring the high consumption strategy. We will refer to population 1 as responsible, and population 2 as irresponsible. The set of all feasible policies specified by Assumptions 2 and 3 is visually depicted in Figure 2.
@@ III Stability analysis of two-population game
1) Suppose $\alpha_{2}>\theta_{1}$. Then a tragedy of the commons is globally asymptotically stable, i.e. $\lim_{t\rightarrow\infty}n(t)=0$.
2) Suppose $\alpha_{2}\leq\theta_{1}$.
@@
    If $\frac{\alpha_{2}-\theta_{1}}{\alpha_{1}+\alpha_{2}}\delta_{SP0}^{1}\leq\delta_{RT0}^{1}\leq\frac{\delta_{TR1}}{\delta_{PS1}}\delta_{SP0}^{1}$, then the only asymptotically stable fixed point is of the form $(x_{1},0,n)$, where
@@
    If $\delta_{RT0}^{1}<\frac{\alpha_{2}-\theta_{1}}{\alpha_{1}+\alpha_{2}}\delta_{SP0}^{1}$, then the only asymptotically stable fixed point constitutes a tragedy of the commons.
Several remarks are in order. In item 1 above, the irresponsible population induces a tragedy of the commons if its consumption rate is higher than the responsible population’s restoration rate. Item 2a provides a region of sustainable policies for population 1. We note that this region gets smaller as $\alpha_{2}$ increases while remaining less than $\theta_{1}$.
@@
Item 2b provides a region where the population 1 policy fails to sustain the resource even though $\alpha_{2}<\theta_{1}$.
@@ IV Results: exploitation of resources
In this section, we study the following hierarchical decision problem posed informally as: how much consumption can the irresponsible population get away with? To approach this question, we consider a local authority for population 2 that may set the consumption rate $\alpha_{2}\geq 0$. This decision is representative of, for example, water usage or fishing regulations.
From Theorem 3.1, we succinctly summarize the resource level that results in the asymptotically stable state for any choice of $\alpha_{2}$ by defining the resource function
@@
where the function $n^{*}(\alpha_{2}):=-\frac{g_{1}(\frac{\alpha_{1}+\alpha_{2}}{\alpha_{1}+\theta_{1}},0)}{\frac{\partial g_{1}}{\partial n}(\frac{\alpha_{1}+\alpha_{2}}{\alpha_{1}+\theta_{1}})}$ is taken directly from (8). The utility that the authority seeks to maximize is defined by
$U(\alpha_{2}):=\alpha_{2}\cdot R(\alpha_{2}).$ (10)
The choice to increase consumption comes at the cost of worsening or even destroying the environmental resource. Thus, (10) captures the tension between resource consumption and the stability of the resource. The optimal consumption rate can be determined by solving the optimization problem
@@
    If $C(\delta_{SP0}^{1})\leq\delta_{RT0}^{1}\leq\frac{\delta_{TR1}^{1}}{\delta_{PS1}^{1}}\delta_{SP0}^{1}$, then
@@
Item a) specifies the range of population 1 environmental policies where population 2 benefits most from the maximal consumption rate $\alpha_{2}^{*}=\theta$. Any higher consumption rate will cause the resource to collapse. Interestingly, in this range, the resource (and consequently, utility) is highly sensitive at the threshold, since $R(\theta_{1})>0$ and discontinuously drops to zero for $\alpha_{2}>\theta_{1}$.
@@ VI Conclusion and Future Work
In this paper, we studied a feedback-evolving game with two populations that share a common environmental resource. This marks initial steps in extending the framework to multi-population interactions. We focused on a particular scenario where one population is “responsible” about using environmental resources, while the other population is “irresponsible”. We characterized the asymptotic dynamic outcomes. We then evaluated to what extent the irresponsible population can take advantage of the resources by deriving optimal consumption rates. Lastly, a sensitivity analysis was provided that suggests incentivizing mutual cooperation is more effective in promoting resource levels under the two-population setting.
This work sets the stage for future studies involving hierarchical decision-making with multiple populations. For example, how would multiple irresponsible populations interact with one another? One can also consider scenarios where populations have their own local environmental resources that have externalities on others.

### v2

@@ I Introduction
Many game-theoretic analyses describe strategic interactions between individuals provided that the incentives for their choices are static, i.e. do not change over time. However, choices often have an impact on a shared environment, which in turn affect their incentives for future choices. The utilization of common resources such as water, fishing grounds, or traffic networks best illustrates this – high individual utilization makes fewer resources available to others in the future [1].
@@
Much of the existing research focuses on a single population of agents whose decisions impact the environment in isolation [8, 9, 10, 7, 11, 12, 13]. These models are not sufficient to describe multi-population interactions. For example, individuals in neighboring countries follow different environmental policies, yet utilize resources from the same common source (e.g. fish in the ocean, water from rivers, clean air). Extensions of feedback-evolving games to multiple populations, and possibly multiple local environments, would enable a vastly richer set of scenarios for study. Recent works in the literature have shifted attention to these multi-population scenarios [14, 15, 16]. For instance, [16] studies epidemic spreading within and between two populations, where the behaviors in each population have externalities on the health state of the other. Overall, an interesting and understudied direction involves hierarchical decision-making, i.e. the strategic selection of local incentive policies given there are environmental externalities between the populations.
@@ II-A Preliminary on Feedback-evolving games
On the other hand, the $A_{0}$ matrix describes payoffs when the environment is depleted. We interpret $A_{0}$ to be an “environmental policy” that the population follows. For example, the low consumption strategy may be more incentivized when the environment is bad (e.g. subsidies for using electric vehicles). The payoff structure of the $A_{0}$ matrix is completely determined from the parameters $\delta_{SP0}:=S_{0}-P_{0}$ and $\delta_{RT0}:=R_{0}-T_{0}$.
@@
We consider initial conditions in the interior $(x,n)\in(0,1)^{2}$, which is forward-invariant under (3). Any initial condition on an edge will stay on that edge. The originating work [2] provided a full characterization of the asymptotic outcomes of system (3) for all possible environmental policy matrices $A_{0}$.
@@
2) Oscillating Tragedy of the commons (OTOC): If $\delta_{SP0}>0$ and $\frac{\delta_{TR1}}{\delta_{PS1}}\delta_{SP0}<\delta_{RT0}$, then system (3) exhibits a stable heteroclinic cycle between the four corner fixed points, $(0,0)$, $(1,0)$, $(1,1)$, $(0,1)$.
3) Tragedy of the commons (TOC): If $\delta_{SP0}\leq 0$, or $\delta_{SP0}>0$ and $\delta_{RT0}<-\frac{\theta}{\alpha}\delta_{SP0}$, then the only asymptotically stable fixed point has $n=0$.
The policies belonging to $\mathcal{V}$ are ones that can sustain a stable and non-zero resource level. In other words, they avert a TOC, which we refer to as a fixed point where the resource level is zero. The policies in $\mathcal{V}$ can be considered as more desirable operating conditions than the policies described in items 2 and 3 above. The OTOC is considered an undesirable outcome, as the system cycles between periods of nearly zero and fully abundant resource levels. Lastly, we note that the time-scale separation $\epsilon$ does not affect the derived stability properties and fixed points of the system.
@@ II-B Model: Two-population feedback-evolving games
where $h(x_{1},x_{2}):=\sum_{i=1}^{2}(\theta_{i}x_{i}-\alpha_{i}(1-x_{i}))$, and with initial condition $(x_{1}(0),x_{2}(0),n(0))\in(0,1)^{3}$. By construction, $(0,1)^{3}$ is forward-invariant. We are interested in studying how the dynamics of the two-population environmental coupled system (7) behaves with respect to the population’s defining parameters – especially how the environmental policies $A_{0}^{(1)}$, $A_{0}^{(2)}$ adopted by each population affects environmental resources. We will make the following assumptions on these environmental policies.
@@
Assumption 2 asserts that the relative payoff to low consumers in both populations monotonically decreases as the environmental state improves. The next and final assumption focuses our study on a scenario where one population 1 is environmentally responsible, and population 2 is irresponsible.
@@
The first part of Assumption 3 asserts that agents in population 1 are cooperative enough to sustain the resource (Theorem 2.1, item 1) in the absence of population 2. The second part asserts that agents in population 2 make no effort to conserve resources, always preferring the high consumption strategy. The set of all feasible policies specified by Assumptions 2 and 3 is visually depicted in Figure 2.
@@ III Stability analysis of two-population game
1) Suppose $\alpha_{2}>\theta_{1}$. Then a tragedy of the commons is asymptotically stable, i.e. $\lim_{t\rightarrow\infty}n(t)=0$.
2) Suppose $\alpha_{2}<\theta_{1}$.
@@
    If $\frac{\alpha_{2}-\theta_{1}}{\alpha_{1}+\alpha_{2}}\delta_{SP0}^{1}<\delta_{RT0}^{1}<\frac{\delta_{TR1}}{\delta_{PS1}}\delta_{SP0}^{1}$, then the only asymptotically stable fixed point is of the form $(x_{1}^{*},0,n^{*})$, where
@@
    If $\delta_{RT0}^{1}\leq\frac{\alpha_{2}-\theta_{1}}{\alpha_{1}+\alpha_{2}}\delta_{SP0}^{1}$, then a tragedy of the commons is the only stable outcome.
3) Suppose $\alpha_{2}=\theta_{1}$. If $\delta_{RT0}^{1}>0$, then there is a locally stable line segment of fixed points given by
$\left\{(1,0,n):n\in\left[0,\frac{\delta_{RT0}^{1}}{\delta_{RT0}^{1}+\delta_{TR1}^{1}}\right)\right\}$ (9)
All other fixed points are isolated and unstable. If $\delta_{RT0}^{1}\leq 0$, then a tragedy of the commons is asymptotically stable.
Several remarks are in order. In item 1 above, the irresponsible population induces a tragedy of the commons if its consumption rate is higher than the responsible population’s restoration rate. Item 2a provides a region of sustainable policies for population 1. This region gets smaller as $\alpha_{2}$ increases while remaining less than $\theta_{1}$. Item 2b provides a region where the population 1 policy fails to sustain the resource even though $\alpha_{2}<\theta_{1}$. Item 3 presents an edge case where there are an infinite number of fixed points, all with different resource levels. The fixed point on the line that the system converges to depends on the initial condition.
@@ IV Results: exploitation of resources
In this section, we study the following hierarchical decision problem posed informally as: how much consumption can the irresponsible population get away with? To approach this question, we consider a local authority for population 2 that may set the consumption rate $\alpha_{2}\geq 0$. This decision is representative of, for example, water usage or fishing regulations. The utility that the authority seeks to maximize is defined by
$U(\alpha_{2}):=\alpha_{2}\cdot R(\alpha_{2}).$ (10)
where $R(\cdot)$ summarizes the resource level that results in the asymptotic outcome of the dynamics. From Theorem 3.1, we define it as
@@
where $n^{*}(\alpha_{2}):=-\frac{g_{1}(\frac{\alpha_{1}+\alpha_{2}}{\alpha_{1}+\theta_{1}},0)}{\frac{\partial g_{1}}{\partial n}(\frac{\alpha_{1}+\alpha_{2}}{\alpha_{1}+\theta_{1}})}$ is taken directly from (8). In the edge case $\alpha_{2}=\theta_{1}$, Theorem 3.1 states there is a range $n\in[0,\frac{\delta_{RT0}^{1}}{\delta_{RT0}^{1}+\delta_{TR1}^{1}}]$ of possible asymptotic resource levels depending on the initial condition. In our definition of $R(\cdot)$, we have elected to assign the highest possible resource [^1] This is done for two reasons. First, it makes the resource function well-defined and left-continuous at $\theta_{1}$. Thus, $R$ attains a maximum value in the interval $[0,\theta_{1}]$. Second, since no particular initial condition is prescribed in the model, we may view the irresponsible population to be “optimistic” regarding the best case among all possible outcomes. value $R(\alpha_{2})=\frac{\delta_{RT0}^{1}}{\delta_{RT0}^{1}+\delta_{TR1}^{1}}$.
The choice to increase consumption comes at the cost of worsening or even destroying the environmental resource. Thus, (10) captures the tension between resource consumption and the stability of the resource. The optimal consumption rate can be determined by solving the problem
@@
    If $C(\delta_{SP0}^{1})\leq\delta_{RT0}^{1}<\frac{\delta_{TR1}^{1}}{\delta_{PS1}^{1}}\delta_{SP0}^{1}$, then
@@
Item a) specifies the range of population 1 environmental policies where population 2 benefits most from the maximal consumption rate $\alpha_{2}^{*}=\theta_{1}$. Any higher consumption rate will cause the resource to collapse. In this range, the resource (and consequently, utility) is highly sensitive at the threshold, since $R(\theta_{1})>0$ and discontinuously drops to zero for $\alpha_{2}>\theta_{1}$.
@@ VI Conclusion and Future Work
In this paper, we studied a feedback-evolving game with two populations that share a common environmental resource. This marks initial steps in extending the framework to multi-population interactions and hierarchical decsion-making. We focused on a scenario where one population is “responsible” about using environmental resources, while the other population is “irresponsible”. We evaluated to what extent the irresponsible population can take advantage of the resources by deriving optimal consumption rates. Lastly, a sensitivity analysis was provided that suggests incentivizing mutual cooperation is more effective than unilateral cooperation. Future work will analyze how multiple irresponsible populations interact with one another. Beyond the scope of the present study, one future direction is to investigate populations that each have local environmental resources. Another interesting consideration is to study evolutionary dynamics beyond the replicator equation.